OpenStudy (anonymous):

Jarod is having a problem with rabbits getting into his vegetable garden, so he decides to fence it in. The length of the garden is 9 feet more than 6 times the width. He needs 74 feet of fencing to do the job. How many feet is the length of the garden?

OpenStudy (anonymous):

assume length of the garden = x length of the garden = y x = 9+ 6y can u do it now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Area = 74 = x*y so x = 9+ 6y x*y = 74 find x now?

OpenStudy (anonymous):

No?